OpenStudy (anonymous):

factor x squared-51x+144?

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} x^2&-51x&+144\\ &\uparrow &\uparrow \\ &-3-16&-3\cdot -16 \end{array}\)

OpenStudy (jdoe0001):

hmmm shoot.... even that's wrong... one sec

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} x^2&-51x&+144\\ &\uparrow &\uparrow \\ &-3-48&-3\cdot -48 \end{array}\)

OpenStudy (anonymous):

so wats the final answer

OpenStudy (jdoe0001):

well... check your material on factoring quadratics

OpenStudy (anonymous):

wat just tell me the answer plz